OpenStudy (tacomaco90):

Experimentation in Physics Help? I cannot understand this question can someone help? Linda wants to measure the effects of mass on the period of a pendulum swing. What Variable should be controlled in her experiment? A Mass of the bob B Length of the string C Period of the swing Wouldnt it be the mass of the bob because she wants to test the different effects of mass on the pendulum?

OpenStudy (astrophysics):

Lets recall the equation of a simple pendulum \[T = 2 \pi \sqrt{\frac{ L }{ g }}\] from all of your above options, what does the period rely on?

OpenStudy (anonymous):

No, that's not the point. That's the equation for an idealized pendulum, but there are a number of reason why any real life pendulum *does* have a mass-dependent oscillation period. I assume by "controlled" they mean fixed, in which case it should be the length of the pendulum. While controlling the length, you would vary the mass.

OpenStudy (astrophysics):

Yes, that's what I was getting at, I wanted the user to understand that the period relies on the length for a simple pendulum, to get the idea. But hey, thanks for checking, so the reason the length is controlled is to show that the change in the period is due to the mass and not to the length of the string.

OpenStudy (masumanwar):

The compound pendulum is in simple harmonic motion when it is in motion it oscillates and the motion must have kinetic and potential energy and the motion dissipates if there is air friction, the motion is in under gravitational acceleration due to gravity .